OpenStudy (anonymous):

Please help.. D: 2. Simplify the following problem using the distributive property: 2x + 3y + 2(x – y) – 3x. Show each step in your calculation. Justify each of your steps using number properties.

OpenStudy (anonymous):

\[2x+3y+2(x-y)-3x=2x+3y+2x-2y-3x=(2x+2x-3x)+(3y-3y)=x+y\]

OpenStudy (anonymous):

\[x+y\]